Please help me solve this equation:
{sqrt( 3x + 7 ) + sqrt( x + 2 )} = 1

Answer by stanbon(75887)

sqrt( 3x + 7 ) + sqrt( x + 2 ) = 1
-------
Square both sides to get:
(3x+7) +2sqrt[3x^2 + 13x + 14] + x+2 = 1
---------
2sqrt[3x^2 + 13x + 14] = -(4x+8)
---------
Square both sides to get:
4(3x^2+13x+14) = 16x^2 + 64x + 64
-----
12x^2 + 52x + 56 = 16x^2 + 64x + 64
-----
4x^2 +12x + 8 = 0
--------
x^2 + 3x + 2 = 0
----
Factor::
(x+2)(x+1) = 0
x = -1 or x = -2
-----
Check those answers in the original equation
to see if either or both are extraneous.
============
Cheers,
Stan H.
====================